# FNOPE: Simulation-based inference on function spaces with Fourier Neural Operators

### v1

@@ 4.3 SIRD: Inference on unseen, non-uniform discretizations
Next, we consider the Susceptible-Infected-Recovered-Deceased (SIRD) model [32] to demonstrate the ability of FNOPE to solve inference problems on non-uniform discretizations of the parameters and observations that were not seen in the training data. In addition, we also show its ability to simultaneously infer vector-valued parameters. The model has three parameters: recovery rate, death rate, and contact rate [33, 34]. We use the same setup as in Gloeckler et al. [30], where we assume that the contact rate varies over time, but recovery and death rates are constant in time. We sample training simulations on a dense uniform grid for both parameters and observations. For evaluation we sample 100 observations, each discretized on a different set of 40 randomly sampled time points in $[0,50]$, using contact rates defined on a distinct set of 40 randomly sampled times.
@@
FNOPE, as well as Simformer, can reliably infer the posterior distribution (Fig. 4a) and the observations lie close to the mean of the posterior predictive (Fig. 4b). Both methods are comparable in terms of MSE of posterior predictive samples to the observations, as well as producing well-calibrated posteriors (Fig. 4c). When we use only 20 timepoints to condition on, the performance of FNOPE slightly decreases (Fig. S1). This highlights the necessity of the FNO block to have enough observation points to perform a reliable (approx.) Fourier transformation. This experiment shows that FNOPE is on par with the state of the art on this low dimensional problem: It successfully infers function-valued parameters together with vector-valued parameters and can be conditioned on arbitrary discretizations of the observations. However, FNOPE can also be applied to very high dimensional problems, as shown in the following experiment.
@@ 4.4 Darcy flow: Scalable inference in high dimensions
The Darcy flow is defined by a second order elliptic PDE and has been used to model many processes including the deformation of linearly elastic materials, or the electric potential in conductive materials. In the geosciences, the Darcy equation is used to describe the distribution of groundwater as a function of the spatially variable hydraulic permeability, which can be inferred from point observations in wells [35]. The Darcy flow is a common benchmark model for FNO applications, especially in the context of training PDE emulator models [13, 36, 37].
@@
where $a(x)\geq 0$ is the permeability we want to infer and $u(x)$ is the hydraulic potential. We adapt the implementation from [38], which provides a GPU-optimized solver. We use a log-normal prior distribution for the permeability, similar to Lim et al. [37]: $b=\log(a)\sim N(0,(-\Delta+\tau I)^{-2})$, where $\Delta$ is the Laplacian operator and $\tau=9$. We sample the prior on a $129\times 129$ grid which results in $\approx 16$k parameters. For FNOPE, we use the first $32$ Fourier modes in both spatial dimensions and for spectral NPE/FMPE we used the first $16$ modes, resulting in $5122\cdot 16^{2}=$ 512 parameter dimensions. For all methods, we infer the log-permeability and evaluate in the original space (as in [37]).
Samples from the posterior inferred with FNOPE closely resemble the ground truth (Fig. 5a). Both FNOPE and FNOPE (fix) correctly capture the fine-structure of the posterior samples and reproduce parameters at much higher fidelity than all baseline methods. While the spectral methods learn oversmoothed posteriors that do not capture local structures, the posterior samples from FMPE (raw) are much noisier and only capture the rough global structure. The posterior means show a similar trend (Fig. S3), and the standard deviations of the baseline methods are higher compared to FNOPE (Fig. S4).
The MSEs between posterior predictive samples and ground truth observations of FNOPE and FNOPE (fix) are consistently better compared to the spectral baseline methods, especially at lower simulation budgets, and are in the same range as FMPE (raw) (Fig. 5b). While all methods are reasonably well-calibrated (Fig. 5c), the visual appearance is vastly different. We additionally measure the posterior quality in terms of posterior log-probability (normalized by the number of pixels) of the associated ground truth parameter $\theta$ [31]. FNOPE has a much higher log probability (Fig. 5d) compared to FMPE (raw). FNOPE (fix) also achieves strong performance for a sufficient number of simulations. As the spectral methods do not model the parameters directly, we cannot calculate the log-probabilities they assign to the ground truth parameters. Overall, FNOPE is the only method that consistently performs well on all presented metrics.
@@ 4.5 Mass balance rates of Antarctic ice shelves: Real world application
Figure 6: Mass balance rates of ice shelves. (a) Measurement transect of the radar data in Ekström Ice Shelf, Antarctica (adapted from [39], published under Creative Commons CC BY license). (b) Posterior predictive results obtained with FNOPE (trained on $10^{5}$ simulations), compared to radar-based observation on one layer in the ice shelf. (c)-(e) Performance measures on test simulations and the real observation, where NPE (raw) refers to the method used in [39].
Finally, we turn to a real-world task from glaciology: Inference of snow accumulation and basal melt rates of Antarctic ice shelves from radar internal reflection horizons (IRHs) [40, 41, 39]. Snow continuously accumulates on top of the ice shelf. Over time, it is transported to larger depths where the former surfaces are further deformed by ice flow and form internal layers of constant age, which are measured by radar (Fig. 6a,b). The inference of accumulation and melt rates is a challenging SBI task, where the model is misspecified, as it cannot accout for all real-world effects. We use an isochronal advection scheme forward model as described in [39]. In this work, the authors consider simulations on a grid of 500 points along a one-dimensional spatial domain and directly infer 50 parameters on a fixed downsampling of this domain using NPE. We refer to this approach as NPE (raw) and compare it to FNOPE and the baseline methods.
First, we evaluate all methods on a test set of simulations (Fig. 6c). As with the previous tasks, the performance of FNOPE at $10^{3}$ simulations is comparable to the performance of the other methods at $10^{5}$ simulations in terms of predictive MSE. All methods show a reasonable calibration in terms of SBC EoD at all simulation budgets (Fig. 6d). We then test the performance of all methods on real data (as in [39]). Posterior predictive samples from FNOPE match the observation very well (Fig.6b), and while FNOPE still performs better at low simulation budgets than the other methods, the relative improvement compared to the baselines is smaller than the one observed on synthetic data (Fig. 6e). We note that this modeling problem was explicitly set up by Moss et al. [39] so that NPE (raw) can infer the posterior using a feasible number of simulations. FNOPE achieves the same performance with two orders of magnitude fewer simulations. Additionally, FNOPE is able to infer the full parameter dimensionality (500) instead of downsampling to 50 dimensions (Fig. S5).
@@ 5 Discussion
Scaling SBI methods to high-dimensional parameter spaces has been the focus of many works which make use of state of the art generative modeling techniques such as generative adversarial networks [7], diffusion models [8, 11, 10] and flow matching [9]. In particular, recent works [30, 42] use a transformer architecture to tokenize function-valued parameters, allowing for complete flexibility in estimating conditional distributions. However, as these methods explicitly model each point for discretized function-valued parameters, they are limited in terms of scalability. Our FNO-based approach allows us to compactly represent the parameters, significantly lowering the computational costs as the number of parameters grows.
Fourier Neural Operators have been previously used for estimating function-valued parameters [43, 44]. Lingsch et al. [45] use a FNO architecture with an FMPE objective to learn vector-valued parameters and, additionally, learn an emulator producing function-valued observations. The crucial difference to our approach is that their simulators are deterministic and the inferred parameters are not function-valued. Recently, Lim et al. [37] developed an approach for score-based modeling in function spaces using FNOs. This enables high-dimensional posterior inference, but their approach is limited to uniform grids and does not allow for flexible conditioning.
Another related approach is diffusion posterior sampling [46, 47, 48], which seeks to learn a high-dimensional prior distribution from samples using score-based models. The learned priors can then be used to generate samples from the posterior using analytically tractable model likelihoods [49, 50]. Other works extended such approaches for intractable likelihoods [51, 52]. Similarly to diffusion posterior sampling, we only require prior samples. However, instead of learning the prior distribution, we learn the posterior distribution directly.
@@
The FNO-backbone used by FNOPE  inherently makes assumptions about the structure of the parameters. These assumptions enabled computationally efficient inference, but result in some limitations. First, the FNO assumes limited high-frequency information in the parameter and observation domains. Therefore, they are ill-suited to infer parameters with high power in higher frequencies—for example, parameters with discontinuities. This could potentially be addressed by neural operators using other transforms, such as wavelet transforms [53]. Furthermore, to (accurately) compute the FFT or NUDFT of the observations, we require sufficiently many points in their discretization. Therefore, unlike other flexible methods [30, 42], our approach cannot perform inference using extremely sparse observations. Still, the SIRD experiment shows that even for 20 points we get reasonable estimations (Fig. S1). Finally, the computational complexity of our approaches still scales exponentially with the dimension of the parameter domain which could be challenging in high-dimensional domains.
@@ Conclusion
This work was funded by the German Research Foundation (DFG) under Germany’s Excellence Strategy – EXC number 2064/1 – 390727645 and SFB 1233 ’Robust Vision’ (276693517) and DFG (DR 822/3-1). This work was co-funded by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A, and the Heinrich-Böll-Stiftung. Data collection was supported by Alfred Wegener Institute through logistic grants AWI_ANT_18. The authors acknowledge support by the state of Baden-Württemberg through bwHPC. GM is a member of the International Max Planck Research School for Intelligent Systems (IMPRS-IS). We thank Manuel Gloeckler for providing the training data and Simformer results for the SIRD experiment. We thank Daniel Gedon and Julius Vetter, and all members of Mackelab for discussions and feedback on the manuscript.
@@
For all baseline SBI methods, we use the sbi toolbox [29], for the Simformer baseline we use the publicly available code from Gloeckler et al. [30] . We use an optimized solver to solve the Darcy Flow PDE [38].
We use various compute resources for the different experiments. For each experiment, we run the training and evaluation for each method, for each simulation budget, and for each of the three random seeds separately. We provide information on the maximum wall-clock time of these runs for each experiment. Code to use FNOPE and reproduce the results is available at https://github.com/mackelab/fnope.
For the Linear Gaussian and SIRD experiments, we perform our experiments on Nvidia RTX 2080ti GPU nodes. Both simulators have negligible wall-clock costs on these GPU nodes. All training runs finished within 2 hours of wall-clock time for all methods and all simulation budgets.
The Darcy flow experiment required GPUs with higher VRAM to accommodate the large ($\approx 16k$ dimensional) parameters and observations. We performed these experiments on Nvidia A100 GPUs for both training and evaluation. Each training run for each method at all budgets was complete within 3 hours of wall-clock time.
For the mass balance rates experiment, we perform training and evaluation on CPU, namely Intel Xeon Gold 16 cores, 2.9GHz. We perform this experiment on CPU as the main computation cost is in running the simulations for the predictive MSE check, and the implementation of the model is not accelerated by use of GPUs. The exact simulation costs are described in Moss et al. [39]. We perform 1000 simulations to calculate the mean predictive MSE in each training run. Each training and evaluation run for all method was completed within 24 hours of wall-clock time.
@@
We here describe more details about our evaluation procedures. We evaluate on a heldout test set $\{(\theta_{j}^{o},l^{\theta}_{j},\eta_{j}^{o},x_{j}^{o},l^{x}_{j})\}_{j=1}^{N_{\text{test}}}$, where $N_{\text{test}}$ is the number of test simulations. Given an approximate posterior distribution $q^{\phi}_{l^{\theta}}(\theta,\eta\mid x,l^{x})$ and a test observation $(x_{j}^{o},l^{x}_{j})$, we draw $N_{\text{post}}$ posterior samples $(\theta_{ij},\eta_{ij})\sim q^{\phi}_{l^{\theta}}(\theta,\eta\mid x^{o}_{j},l^{x}_{j})$. In the case where no vector-valued parameters $\eta$ are present, they can be omitted. Similarly, for methods which do not explicitly use the positions $l^{\theta},l^{x}$ (e.g. FNOPE  (fix)), the positions can be omitted, as we do not apply these methods to tasks where we consider arbitrary discretizations.
We report the average and standard error over all $N_{\text{test}}$ test simulations.
@@
We calculate the sliced Wassertein(-2) distance (SWD) [27], with 50 random projections.
@@
Simulation-based calibration (SBC) [28] is a standard measure of the calibration of approximate posterior distributions (in terms of over- or underconfidence). We obtain ranks $r_{ij}$ for each sample $(\theta_{j}^{o},x_{j}^{o})$ in the test set using SBC with the 1-dimensional marginal distributions used as the reducing functions. That is, for each of the dimensions $i$ of $\theta$, the rank $r_{ij}$ is an integer in $(1,N_{\text{post}}+1)$. This results in $N_{\text{test}}\times d_{\theta}$ ranks. The cumulative distribution function of ranks is therefore
$\text{CDF}(\alpha)=\frac{1}{N_{\text{test}}\cdot d_{\theta}}\sum_{j}\mathbb{I}[r_{j}/N_{\text{post}}<\alpha].$
@@
In contrast to the SBC area under the curve (SBC AUC), the EoD will detect poor calibrations for posteriors that are overconfident at low confidence levels $\alpha$, and underconfident at high $\alpha$ (or vice-versa).
S2.3 Predictive MSE
To calculate the MSE for posterior predictive samples, we run for each posterior sample $(\theta_{ij},\eta_{ij})$, and each true observation $x_{j}^{o}$, the simulator $x_{ij}\sim p_{l^{x}_{j}}(x\mid\theta_{ij},l^{\theta},\eta_{ij})$. We then compute the average mean square error of the simulation $x_{ij}$ to the corresponding observations $x_{j}^{o}$,
$\text{MSE}=\frac{1}{N_{\text{test}}N_{\text{post}}}\sum_{i=1,j=1}^{N_{\text{test}},N_{\text{post}}}\frac{1}{|l^{x}_{j}|}||x_{ij}-x_{j}^{o}||_{L^{2}}^{2},$
@@
For the Darcy Flow task, we additionally report the posterior log-probability of the true parameters [31], normalized by the number of pixels:
[…]
@@
S3.1 Kernel lengthscale heuristic
@@
[…]
We consider discretizations normalized to $[0,1]$ in each dimension, and so the power contained in the first $M$ spectral modes, $\bar{P}_{M}$, corresponds to the above integral within the domain $||f||_{\infty}\leq M/2$, i.e. where all the components of $f$ are within $[-M/2,M/2]$. Therefore $\bar{P}_{M}$ simplifies to the product of the Gaussian integrals
[…]

### v2

@@ 4.3 SIRD: Inference on unseen, non-uniform discretizations
Next, we consider the Susceptible-Infected-Recovered-Deceased (SIRD) model [kermack1927contribution_sir] to demonstrate the ability of FNOPE to solve inference problems on non-uniform discretizations of the parameters and observations that were not seen in the training data. In addition, we also show its ability to simultaneously infer vector-valued parameters. The model has three parameters: recovery rate, death rate, and contact rate [chen2020timeSIR, schmidt2021probabilistic]. We use the same setup as in glockler2024simformer, where we assume that the contact rate varies over time, but recovery and death rates are constant in time. We sample training simulations on a dense uniform grid for both parameters and observations. For evaluation we sample 100 observations, each discretized on a different set of 40 randomly sampled time points in $[0,50]$, using contact rates defined on a distinct set of 40 randomly sampled times.
@@
FNOPE, as well as Simformer, can reliably infer the posterior distribution (Fig. 4a) and the observations lie close to the mean of the posterior predictive (Fig. 4b). Both methods are comparable in terms of MSE of posterior predictive samples to the observations, as well as producing well-calibrated posteriors (Fig. 4c). When we use only 20 timepoints to condition on, the performance of FNOPE slightly decreases (Fig. S2). This highlights the necessity of the FNO block to have enough observation points to perform a reliable (approx.) Fourier transformation. We also observe that FNOPE performs robustly across the base distribution lengthscale (Fig. S3). This experiment shows that FNOPE is on par with the state of the art on this low dimensional problem: It successfully infers function-valued parameters together with vector-valued parameters and can be conditioned on arbitrary discretizations of the observations. However, FNOPE can also be applied to very high dimensional problems, as shown in the following experiment.
@@ 4.4 Darcy flow: Scalable inference in high dimensions
The Darcy flow is defined by a second order elliptic PDE and has been used to model many processes including the deformation of linearly elastic materials, or the electric potential in conductive materials. In the geosciences, the Darcy equation is used to describe the distribution of groundwater as a function of the spatially variable hydraulic permeability, which can be inferred from point observations in wells [NowakHydrology]. The Darcy flow is a common benchmark model for FNO applications, especially in the context of training PDE emulator models [li2021FNO, li2024physics, lim2025DDO].
@@
where $a(x)\geq 0$ is the permeability we want to infer and $u(x)$ is the hydraulic potential. We adapt the implementation from [nemo2023], which provides a GPU-optimized solver. We use a log-normal prior distribution for the permeability, similar to lim2025DDO: $b=\log(a)\sim N(0,(-\Delta+\tau I)^{-2})$, where $\Delta$ is the Laplacian operator and $\tau=9$. We sample the prior on a $129\times 129$ grid which results in $\approx 16$k parameters. For FNOPE, we use the first $32$ Fourier modes in both spatial dimensions and for spectral NPE/FMPE we used the first $16$ modes, resulting in $2\cdot 16^{2}=512$ parameter dimensions. For all methods, we infer the log-permeability and evaluate in the original space (as in [lim2025DDO]).
Samples from the posterior inferred with FNOPE closely resemble the ground truth (Fig. 5a and Fig. S6). Both FNOPE and FNOPE (fix) correctly capture the fine-structure of the posterior samples and reproduce parameters at much higher fidelity than all baseline methods. While the spectral methods learn oversmoothed posteriors that do not capture local structures, the posterior samples from FMPE (raw) are much noisier and only capture the rough global structure. The posterior means show a similar trend (Fig. S7), and the standard deviations of the baseline methods are higher compared to FNOPE (Fig. S8).
The MSEs between posterior predictive samples and ground truth observations of FNOPE and FNOPE (fix) are consistently better compared to the spectral baseline methods, especially at lower simulation budgets, and are in the same range as FMPE (raw) (Fig. 5b). While all methods are reasonably well-calibrated (Fig. 5c), the visual appearance is vastly different. We additionally measure the posterior quality in terms of posterior log-probability (normalized by the number of pixels) of the associated ground truth parameter $\theta$ [lueckmann2021benchmarking]. FNOPE has a much higher log probability (Fig. 5d) compared to FMPE (raw). FNOPE (fix) also achieves strong performance for a sufficient number of simulations. As the spectral methods do not model the parameters directly, we cannot calculate the log-probabilities they assign to the ground truth parameters. Overall, FNOPE is the only method that consistently performs well on all presented metrics. In additional ablation experiments (Fig. S4), we show that FNOPE attains strong performance across different hyperparameter choices, and that our lengthscale heuristic (Sec. 3.1) is an appropriate choice for this task.
@@ 4.5 Mass balance rates of Antarctic ice shelves: Real world application
Figure 6: Mass balance rates of ice shelves. (a) Measurement transect of the radar data in Ekström Ice Shelf, Antarctica (adapted from [moss2023simulation], published under Creative Commons $C$ BY license). (b) Posterior predictive results obtained with FNOPE (trained on $10^{5}$ simulations), compared to radar-based observation on one layer in the ice shelf. (c)-(e) Performance measures on test simulations and the real observation, where NPE (raw) refers to the method used in [moss2023simulation].
Finally, we turn to a real-world task from glaciology: Inference of snow accumulation and basal melt rates of Antarctic ice shelves from radar internal reflection horizons (IRHs) [Waddington2007, Wolovick2021, moss2023simulation]. Snow continuously accumulates on top of the ice shelf. Over time, it is transported to larger depths where the former surfaces are further deformed by ice flow and form internal layers of constant age, which are measured by radar (Fig. 6a,b). The inference of accumulation and melt rates is a challenging SBI task, where the model is misspecified, as it cannot accout for all real-world effects. We use an isochronal advection scheme forward model as described in [moss2023simulation]. In this work, the authors consider simulations on a grid of 500 points along a one-dimensional spatial domain and directly infer 50 parameters on a fixed downsampling of this domain using NPE. We refer to this approach as NPE (raw) and compare it to FNOPE and the baseline methods.
First, we evaluate all methods on a test set of simulations (Fig. 6c). As with the previous tasks, the performance of FNOPE at $10^{3}$ simulations is comparable to the performance of the other methods at $10^{5}$ simulations in terms of predictive MSE, and is only marginally worse at $10^{2}$ simulations. All methods show a reasonable calibration in terms of SBC EoD at all simulation budgets (Fig. 6d). We then test the performance of all methods on real data (as in [moss2023simulation]). Posterior predictive samples from FNOPE match the observation very well (Fig.6b), and while FNOPE still performs better at low simulation budgets than the other methods, the relative improvement compared to the baselines is smaller than the one observed on synthetic data (Fig. 6e). We note that this modeling problem was explicitly set up by moss2023simulation so that NPE (raw) can infer the posterior using a feasible number of simulations. FNOPE achieves the same performance with two orders of magnitude fewer simulations. Additionally, FNOPE is able to infer the full parameter dimensionality (500) instead of downsampling to 50 dimensions (Fig. S5).
@@ 5 Discussion
Scaling SBI methods to high-dimensional parameter spaces has been the focus of many works which make use of state of the art generative modeling techniques such as generative adversarial networks [ramesh2022gatsbi], diffusion models [geffner2022npse, linhart2024diffusionposterior, schmitt2024consistency] and flow matching [wildberger2023fmpe]. In particular, recent works [glockler2024simformer, Chang2024AmortizedConditioningEngine] use a transformer architecture to tokenize function-valued parameters, allowing for complete flexibility in estimating conditional distributions. However, as these methods explicitly model each point for discretized function-valued parameters, they are limited in terms of scalability. Our FNO-based approach allows us to compactly represent the parameters, significantly lowering the computational costs as the number of parameters grows. Finally, for applications where only the one-dimensional marginal distributions of the Bayesian posterior are needed, it is possible to scale SBI methods to higher dimensions, as the correlation between the parameters does not need to be captured [ambrogioni2019favi]. However, in most scientific applications the correlation structure is an essential object of interest and marginal distributions are only of limited use.
Estimating function-valued parameters using FNOs [GANO, VANO, behroozi2025SCFNO] and other neural operators [Hao2023GNOT, Wang2024LNO, wu2024Transolver] has been explored in previous work. A majority of these approaches consider deterministic inversion, or estimating a single value for the parameters as opposed to targeting the Bayesian posterior. While probabilistic generative models such as invertible Fourier Neural Operators (iFNOs, long2025iFNO) estimate a conditional distribution, they do not explicitly target the Bayesian posterior. We compare the performance of iFNO on the Darcy task, and while it has comparable performance in terms of predictive MSE, the uncertainty estimates are not well-calibrated and the conditional distribution collapses to a tiny parameter region (Appendix S8).
The closest neighbours to our work are lingsch2024fuse, who use a FNO architecture with an FMPE objective to learn vector-valued parameters and, additionally, learn an emulator producing function-valued observations. The crucial difference to our approach is that their simulators are deterministic and the inferred parameters are not function-valued. Recently, lim2025DDO developed an approach for score-based modeling in function spaces using FNOs. This enables high-dimensional posterior inference, but their approach is limited to uniform grids and does not allow for flexible conditioning.
Another related approach is diffusion posterior sampling [song2021scorebased, song2023pseudoinverseguided, chung2023diffusion], which seeks to learn a high-dimensional prior distribution from samples using score-based models. The learned priors can then be used to generate samples from the posterior using analytically tractable model likelihoods [wang2023zeroshot, kawar2022restoration]. Other works extended such approaches for intractable likelihoods [cardoso2024monte, wu2024principled]. Similarly to diffusion posterior sampling, we only require prior samples. However, instead of learning the prior distribution, we learn the posterior distribution directly.
@@
The FNO-backbone used by FNOPE  inherently makes assumptions about the structure of the parameters. These assumptions enable computationally efficient inference, but result in some limitations. First, the FNO assumes limited high-frequency information in the parameter and observation domains. Therefore, they are ill-suited to infer parameters with high power in higher frequencies—for example, parameters with discontinuities. This could potentially be addressed by neural operators using other transforms, such as wavelet transforms [tripura2023wavelet]. Furthermore, to (accurately) compute the FFT or NUDFT of the observations, we require sufficiently many points in their discretization. Therefore, unlike other flexible methods [glockler2024simformer, Chang2024AmortizedConditioningEngine], our approach cannot perform inference using extremely sparse observations. Still, the SIRD experiment shows that even for 20 points we get reasonable estimations (Fig. S2). Finally, the computational complexity of our approaches still scales exponentially with the domain dimension, which could be challenging in high-dimensional domains.
@@ Conclusion
This work was funded by the German Research Foundation (DFG) under Germany’s Excellence Strategy – EXC number 2064/1 – 390727645 and SFB 1233 ’Robust Vision’ (276693517) and DFG (DR 822/3-1) and the Heinrich-Böll-Stiftung. Data collection was supported by Alfred Wegener Institute through logistic grants AWI_ANT_18. The authors acknowledge support by the state of Baden-Württemberg through bwHPC. GM is a member of the International Max Planck Research School for Intelligent Systems (IMPRS-IS). We thank Manuel Gloeckler for providing the training data and Simformer results for the SIRD experiment. We thank Daniel Gedon and Julius Vetter, and all members of Mackelab for discussions and feedback on the manuscript.
@@
For all baseline SBI methods, we use the sbi toolbox [BoeltsDeistler_sbi_2025], for the Simformer baseline we use the publicly available code from glockler2024simformer . We use an optimized solver to solve the Darcy Flow PDE [nemo2023].
We use various compute resources for the different experiments. For each experiment, we run the training and evaluation for each method, for each simulation budget, and for each of the three random seeds separately. The summary of network sizes and compute times for all tasks are provided in Tab. S1,S2,S3.
For the Linear Gaussian and SIRD experiments, we perform our experiments on Nvidia RTX 2080ti GPU nodes. Both simulators have negligible wall-clock costs on these GPU nodes. The Darcy flow experiment required GPUs with higher VRAM to accommodate the large ($\approx 16k$ dimensional) parameters and observations. We performed these experiments on Nvidia A100 GPUs. For the Antarctic Ice experiment, we perform training and evaluation on CPU, namely Intel Xeon Gold 16 cores, 2.9GHz. We perform this experiment on CPU as the main computation cost is in running the simulations for the predictive MSE check, and the implementation of the model is not accelerated by use of GPUs. The simulation costs are described in moss2023simulation.
Table S1: Network sizes and training times for Gaussian Linear (GL) and Darcy Flow (Darcy) tasks. We report the mean over 3 runs with a training budget of 10k simulations. Baselines estimating the "raw" (untransformed) parameters are denoted (R) and those estimating Fourier coefficients are denoted (S). Linear Gaussian task was run on a 2080ti GPU, Darcy flow on an A100 GPU.
Task Metric FNOPE FNOPE (fix) NPE (S) FMPE (S) FMPE (R)
GL # params. 110K 109K 567K 86.3K 299K
GL train (Tot.) [m] 3.12 1.98 6.77 2.46 5.14
GL train (/epoch) [s] 0.99 0.87 2.17 0.24 0.31
GL sample (/sample) [ms] 2.16 1.06 0.05 0.24 0.20
Darcy # params. 11.6M 11.6M 3.54M 898K 9.18M
Darcy train (Tot.) [m] 72.2 41.8 20.5 10.9 12.2
Darcy train (/epoch) [s] 20.3 26.2 2.82 0.66 0.73
Darcy sample (/sample) [ms] 280 35.2 0.21 1.95 2.70
Table S2: Network sizes and training times for SIRD. We report the mean over 3 runs with a training budget of 10k simulations. Both methods were trained and evaluated on a 2080ti GPU.
Task Metric FNOPE Simformer
SIRD # params. 117K 286K
SIRD train (Tot.) [m] 6.95 47.17
SIRD train (/epoch) [s] 1.19 9.43
SIRD sample (/sample) [ms] 0.22 0.22
Table S3: Network sizes and training times for Antarctic Ice (Ice) task. We report the mean over 3 runs with a training budget of 10k simulations. Baselines estimating the "raw" (untransformed) parameters are denoted (R) and those estimating Fourier coefficients are denoted (S). All methods were trained and evaluated on CPU.
Task Metric FNOPE NPE (S) FMPE (S) NPE (R) FMPE (R)
Ice # params. 25.3K 236K 92.3K 361K 96.2K
Ice train (Tot.) [m] 47.4 19.7 31.6 6.80 33.5
Ice train (/epoch) [s] 9.82 4.21 1.92 5.64 2.02
Ice sample (/sample) [ms] 22.8 1.50 16.6 1.53 23.9
@@
We here describe more details about our evaluation procedures. We evaluate on a heldout test set $\{(\theta_{j}^{o},l^{\theta}_{j},\eta_{j}^{o},x_{j}^{o},l^{x}_{j})\}_{j=1}^{J_{\text{test}}}$, where $J_{\text{test}}$ is the number of test simulations. Given an approximate posterior distribution $q^{\phi}_{l^{\theta}}(\theta,\eta\mid x,l^{x})$ and a test observation $(x_{j}^{o},l^{x}_{j})$, we draw $K_{\text{post}}$ posterior samples $(\theta_{kj},\eta_{kj})\sim q^{\phi}_{l^{\theta}}(\theta,\eta\mid x^{o}_{j},l^{x}_{j})$. In the case where no vector-valued parameters $\eta$ are present, they can be omitted. Similarly, for methods which do not explicitly use the positions $l^{\theta},l^{x}$ (e.g. FNOPE  (fix)), the positions can be omitted, as we do not apply these methods to tasks where we consider arbitrary discretizations.
We report the average and standard error over all $J_{\text{test}}$ test simulations.
@@
Following bonneel2015slicedwasserstein, we define the (empirical) sliced Wasserstein(-2) distance (SWD) between $N$ samples from two probability distributions $p$ and $q$ as
[…]
where $x_{k}\sim p(x),y_{k}\sim q(y)$ are samples from the two distributions, $u$ are uniformly randomly sampled vectors on the unit sphere $\mathbb{S}^{D-1}$, and $x_{u}^{(k)},y_{u}^{(k)}$ are the 1-dimensional $i$-th order statistics of the projections $u^{\top}x_{k},u^{\top}y_{k}$ respectively. We calculate the SWD with 50 random projections $u$ and $K=1000$ posterior samples.
@@
Simulation-based calibration (SBC) [talts2018validating] is a standard measure of the calibration of approximate posterior distributions (in terms of over- or underconfidence). We obtain ranks $r_{ij}$ for each sample $(\theta_{j}^{o},x_{j}^{o})$ in the test set using SBC with the 1-dimensional marginal distributions used as the reducing functions. That is, for each of the dimensions $i$ of $\theta$, the rank $r_{ij}$ is an integer in $(1,K_{\text{post}}+1)$. This results in $J_{\text{test}}$ ranks. The cumulative distribution function of ranks is therefore
$\text{CDF}_{i}(\alpha)=\frac{1}{J_{\text{test}}}\sum_{j}\mathbb{I}[r_{ij}/K_{\text{post}}<\alpha].$
@@
$\text{SBC EoD}(i)=\int_{0}^{1}|\text{CDF}_{i}(\alpha)-\alpha|d\alpha.$
In contrast to the SBC area under the curve (SBC AUC), the EoD will detect poor calibrations for posteriors that are overconfident at low confidence levels $\alpha$, and underconfident at high $\alpha$ (or vice-versa).Finally, we report the average SBC EoD across the dimensions of $\theta$,
$\text{SBC EoD}=\frac{1}{N_{\theta}}\text{SBC EoD}(i).$
For SIRD, since the posterior dimensionalities are low, we compute an average SBC EoD over all one-dimensional marginals of the posterior. For the Darcy Flow and Antarctic Ice tasks, we select a subset of 50 marginal distribution for computing the SBC EoD, regularly spread across the domain.
S2.3 Predictive MSE
To calculate the MSE for posterior predictive samples, we run for each posterior sample $(\theta_{kj},\eta_{kj})$, and each true observation $x_{j}^{o}$, the simulator $x_{kj}\sim p_{l^{x}_{j}}(x\mid\theta_{kj},l^{\theta},\eta_{kj})$. We then compute the average mean square error of the simulation $x_{kj}$ to the corresponding observations $x_{j}^{o}$,
@@
For the Darcy Flow task, we additionally report the posterior log-probability of the true parameters [lueckmann2021benchmarking], normalized by the number of pixels:
$\text{log-probability per pixel}=\frac{1}{|l^{\theta}_{j}|}\log q^{\phi}_{l^{\theta}_{j}}(\theta^{o}_{j}\mid x^{o}_{j},l^{x}_{j}).$
@@
S3.1 Bochner’s Theorem
We state Bochner’s theorem following williams2006gps. A complex-valued function $k$ on $\mathbb{R}^{d}$ is the covariance function of a weakly stationary mean continuous complex-valued random process on $\mathbb{R}^{D}$ if and only if it can be represented as
$k(\tau)=\int_{\mathbb{R}^{D}}\exp^{2\pi if\cdot\tau}d\mu(s)$ (5)
for some positive finite measure $\mu$. Crucially, in the less general but relevant case that $\mu$ admits a density $P(f)$, the integral is a Fourier transform between the kernel $k(\tau)$ and the spectral density $P(f)$. We apply this result to relate the lengthscale of the square exponential kernel to the spectral density of its Fourier decomposition in Sec. 3.1.
S3.2 Kernel lengthscale heuristic
@@
$\bar{P}=\int_{\mathbb{R}^{d_{\theta}}}P(f)df=1$
We consider discretizations normalized to $[0,1]$ in each dimension, and so the power contained in the first $M$ spectral modes, $\bar{P}_{M}$, corresponds to the above integral within the domain $f||_{\infty}\leq M/2$, i.e. where all the components of $f$ are within [M/2,M/2][-M/2,M/2]. Therefore $\bar{P}_{M}$ simplifies to the product of the Gaussian integrals
$\bar{P}_{M}$ $=\left(\int_{-M/2}^{M/2}(2\pi l^{2})^{1/2}\exp(-2\pi^{2}l^{2}f_{i})df_{i}\right)^{d_{\theta}}$